OpenStudy (anonymous):

You are making poster board boundary markers in the shape of square pyramids for a relay race. If you are including a 10-inch square as the bottom and you want the slant height to be 12 inches, how much poster board do you need?

OpenStudy (anonymous):

They're looking for the surface area of a square pyramid. Base Side = S = 10 Slant Height = L = 12 Base Area = s^2 Triangle Area = (1/2)base*height = (1/2)S*L So the total surface area = Base Area + 4*Triangle Area (since there's 4 sides on a square pyramid). That would be: Surface Area = s^2+4(1/2)SL = S^2+2SL. Solve!

OpenStudy (anonymous):

thank you